ADNAN MUHAMMED

Grade 12,

a body cools from 60 degree celsius to 50 degree celsius in 10 minutes.if the room temperature is 25 degree celsius and assuming newton's law of cooling to hold good,the temperature of the body at the end of the next 10 minutes will be: a)45 deg. cel. b)42.85 deg. cel. c)40 deg. cel. d)38.5 deg cel. (MAHE 2002)


a body cools from 60 degree celsius to 50 degree celsius in 10 minutes.if the room temperature is 25 degree celsius and assuming newton's law of cooling to hold good,the temperature of the body at the end of the next 10 minutes will be: a)45 deg. cel. b)42.85 deg. cel. c)40 deg. cel. d)38.5 deg cel. (MAHE 2002)


Grade:upto college level

3 Answers

Ramreddy IIT Patna
9 years ago

T(t)- T= (To-Ta) e^{-kt}
50-25 = (60-25) e^(-k*10*60)
we can find the value of k from above equation
the temperature of the body at the end of the next 10 minutes can be calculated as below
T-25 = (50-25) e^(-k*10*60)
from above equation we can calculate the value of T.
Amol Singh
7 years ago
T(t)- T= (To-Ta) e^{-kt}50-25 = (60-25) e^(-k*10*60)we can find the value of k from above equationthe temperature of the body at the end of the next 10 minutes can be calculated as belowT-25 = (50-25) e^(-k*10*60)from above equation we can calculate the value of T.
hence found
Mohammed Ehtesham
6 years ago
d\Theta = T dt = 10 \Theta = 3T+2T/2 By formula , d\Theta /dt = K(\Theta - \Theta 0 ) T/10 = K( 3/2T) K = 2/30T Equation(1) For 2nd condition d\Theta = 2T-x dt = 10 mins \Theta = 2T + x / 2 By formula We get 2T-x/10 = K(2T + x / 2 - T) Equation(2) \therefore By solving equations 1 & 2 We get X = 1.5 T